A student found that a brass cylinder weighs 75.04 g in air and 66.11 g in water (density is 0.998 g/cm3). What is the density of brass?


1
Expert's answer
2021-03-12T06:30:22-0500

The space occupied by water ;


(74.04g-66.11g)=7.93g


Volume ="\frac{Mass of displaced water} {Density}" ="\frac{7.93g}{0.998g/cm^{3} }"


Volume =7.9459"cm^{3}"


Volume of Brass ="\frac{Weight of the cylinder} {Volume}"


"=\frac{74.04g}{7.9459cm^{3}}"


"=9.318g\/cm^{3}"